# Unveiling Typographic Deceptions: Insights of the Typographic Vulnerability in Large Vision-Language Model

### v2

@@ 5 Experiments
Evaluation on TypoD: In the beginning, through adopting the data of the Factor Exploring Stage in TypoD, we explore various typographic factors that could affect distractibility, including Font Size, Font Opacity, Spatial Positioning and Font Color for added typographic text in Object Recognition (Obj), Visual Attribute (Vis), Enumeration (Enu) as well as Reasoning (Rea) sub-tasks. In this section, we present the the performance of object recognition, which best represents the VQA task, under various typical typographic factors, as illustrated in Fig. 6 (a). More detailed results with more various factors of other sub-tasks are show in Appendix Sec. 7.
Regarding Font Size and Opacity, Fig. 6 (a) shows a conspicuous positive correlation between the visibility of typographic text and the efficacy of typographic attacks. Furthermore, less visible typographic text still result in a significant reduction in models’ accuracy. typographic text with a font size of 6 pixels and an opacity of 20% can still cause an 11% and 16.6% drop in accuracy, respectively. Concerning Spatial Positioning and Font Color, Fig. 6 (a) shows that typographic text, regardless of their color or position, can significantly impact the accuracy of models. The phenomenon of typographic attacks to remain effective despite variations in color and location may be utilized to craft images laden with typographic text that, while seemingly ordinary, contain inconspicuous errors that can mislead LVLMs. This can be achieved by selecting a font color that merges with the background, making it hard to spot, and positioning it in an area that tends not to draw the eye.
According to the exploration results of typographic factors in Fig. 6 and results in Appendix Sec. 7, we fix the final typographic factors as {15px, 100%, R2C2 and White}. Using this setting, the sub-datasets, namely TypoD-B and TypoD-L, in the Factor Fixing stage of our TypoD are introduced. Next, we validate the performance degradation of LVLMs on Obj, Vis, Enu and Rea in Typo-D and Typo-L. In Typo-B, compared to the performance of LLaVA-v1.5 and InstructBLIP in WTypo-B, there would occur and performances drop. In Typo-L, the performances drop compared to WTypo-L are and respectively. We observe that typographic attacks indeed significantly degrade the performance of LVLMs. However, there are differences in the robustness of different LVLMs models against this threat, with InstructBLIP outperforming LLaVA-v1.5. For a more detailed exploration, we utilized LLaVA-v1.5 in our exploratory phase of LVLMs.
@@
Exploring Discoveries: In order to further validate the discoveries in Fig. 4, we initially examine more image instances in the Appendix Sec. 8 to observe the similarity score in vision-language modal comparison and verify the distract effects of visual attention in Grad-CAM images. In addition, to assess the impact of changes in input information of the text encoder of CLIP on the severity of typographic threats in the zero-shot classification task, we randomly select 500 images from ImageNet and create their typographic versions by printing a random class as a typographic error in images. Hence we get a set of normal images and a set of typographic images. Then we craft image-matching texts for these images using Set 1 and Set 2 templates. Set 1: [an image of {label}, an image of {typo}]. Set 2: [an image of {label} with a word {typo} written on top of it, an image of {typo} with a word {typo} written on top of it, an image of {label}, an image of {typo}]. Through separate testing on CLIP-Base (CLIP-B) and CLIP-Large (CLIP-L), according to the Fig. 6, adopting the more informative Set 2 as templates would result in the enhancements of and compared to the less informative Set 1 in the zero-shot classification task.
For validating the dicoveries in Fig. 5, we need to augment prompts with additional semantics related to ”describing image content” to obtain new responses in text language modality. The text response generated by this new prompt, according to Discovery 2, on one hand, could mitigate the visual attention already attracted by the typographic text due to the increased information in the text modality. Also according to Discovery 3, on the other hand, the additonal information in the text modality can also match attention with the prompt, further directing overall attention favorably towards the correct response. we specifically implement the semantics of the above prompts by: 1. Directly adding requirements for describing images to the prompt; 2. Using multi-step prompts to generate more text language responses. Exemplary instructed prompts are illustrated following.
@@
Compared to the GAP of for LLaVA-v1.5 in TypoD-B as shown in Table 2, the three prompts we test in Table 3 result in reductions in GAP to , , and , respectively. This indicates that adding semantics related to image description in prompts is indeed effective, but there may be a threshold, which may arise from the inherent gap in querying between multi-modal information. Additionally, by comparing the GAP of , , and for Prompt 2.1, 2.2, and 2.3, it further demonstrates that through the incremental increase of semantics describing image content, there is a slight stair-step suppression effect on typographic threats. More results of various prompts are presented in Appendix Sec. 8.
Ablation Study: Except for the above four multi-modal sub-tasks, in the Appendix Sec. 9 we also explore the performance of typographic vulnerability in the textual image. Additionally, we also tested a more intuitive approach to prompt semantic modification for typographic attacks. As presented in Fig. 6 (c), a further GAP decrease could be attained by inserting ”unaffected by typographic text” in the above new Prompts.
@@ 6 Conclusion
In this paper, we conduct verification of typographic attacks on leading LVLMs and confirm their prevalent existence. We presented an extensive Typographic Dataset to delve deeper into this susceptibility, evaluating typographic risks across different multi-modal tasks and typographic factors. Our research reveals three principal findings regarding the effects of typographic attacks on LVLMs. Inspired by these findings, through experimental validation with the Typographic Dataset, we reduce performance degradation from to when LVLMs confront typographic attacks.
Supplementary Material
7 Evaluation on TypoD
7.1 Details of Evaluated LVLMs
@@
7.2 More Quantitative Results
@@
8 Exploring Discoveries
8.1 CLIP with Informative Text Encoder Inputs
CLIP could almost avoid typographic attacks in zero-shot classification if more informative prompts are provided, as demonstrated in Fig. 4. In this subsection, we provide more examples to illustrate above discovery as presented in Fig. 7. By providing additional information to the prompt text, such as including more options like ”an image of a {stingray, tape player} with the word ’tape player’ written on it, CLIP shifts its attention area and correctly distinguishes between the visual contents and the added typographic texts under a typographic attack. This demonstrates that CLIP itself is not flawed and its vision encoder can accurately extract complete information. The crucial issue lies in the varying levels of information contained in the text options provided for the text encoder.
8.2 LLaVA with Informative Prompts
8.3 Performance of Different Informative Prompts in LLaVA-v1.6
For further validating the discoveries in Fig. 5, we evaluate informative prompts in Table 3 on LLaVA-v1.6, as shown in Table 6. Compared to the GAP of for LLaVA-v1.6 in TypoD-B, the five prompts we test result in narrower GAPs of , , , and , respectively. This indicates that adding semantics related to image description in prompts is indeed effective.
9 Ablation Study
9.1 Impacts of Typographic Texts on Images Containing Texts
In order to evaluate how typographic texts affect the performance of LVLMs on images containing texts. We craft a typographic arithmetic dataset. In the arithmetic computation, LVLMs need to comprehend arithmetic expressions embedded in images. Here, we focus on relatively easy arithmetic problems like addition, subtraction, multiplication, and division. We randomly selected two numbers and chose an arithmetic operation to compose an arithmetic problem. The correct answer is assigned as the ground truth, and a random number is selected as a typo. Then both the arithmetic problem and the typographic text are printed into the image. The distractibility of LVLMs with different typographic factors in the arithmetic computation task is shown in Table 12.
9.2 Prompts Instructed to Ignore Typographic Texts
Additionally, we test a more intuitive approach to prompt semantic modification for typographic attacks on LLaVA-v1.5. We add the additional words ”unaffected by typographic texts.” to each Prompts mentioned in Table 3 respectively. Then we get new Prompt 1-1, Prompt 2-1 and Prompt 3-1 which are instructed to ignore typographic texts. As shown in Table 7, compared to GAPs of unmodified prompts in Table 3 which are , and respectively, their instructed versions of ignoring typographic texts result in narrower GAPs of , , . This suggests that incorporating instructions to overlook typographic texts within prompts can genuinely enhance performance.
@@ 9.2 Prompts Instructed to Ignore Typographic Texts
Object Recognition | 98.4 | 49.6 | 48.8 | 97.6 | 77.3 | 20.2 | 98.1 | 78.9 | 19.2 | 97.6 | 81.5 | 16.0 | 97.0 | 84.3 | 12.6 |
Visual Attribute | 92.3 | 82.0 | 10.2 | 97.9 | 94.8 | 3.1 | 97.9 | 95.8 | 2.0 | 96.9 | 95.3 | 1.5 | 97.4 | 95.3 | 2.0 |
Enumeration | 76.0 | 50.7 | 25.2 | 91.8 | 70.2 | 21.5 | 91.8 | 71.3 | 20.5 | 91.5 | 76.5 | 15.0 | 92.8 | 77.6 | 15.2 |
Reasoning | 90.3 | 44.8 | 45.4 | 89.5 | 60.0 | 29.4 | 86.2 | 63.6 | 22.5 | 86.1 | 67.6 | 18.4 | 83.2 | 67.6 | 15.5 |

### v3

@@ 5 Experiments
Evaluation on TypoD: By adopting the data of the Factor Exploring Stage in TypoD, we explore various typographic factors that could affect distractibility, including Font Size, Font Opacity, Spatial Positioning, and Font Color for added typographic text in Object Recognition (Obj), Visual Attribute (Vis), Enumeration (Enu) as well as Reasoning (Rea) sub-tasks. In this section, we present the performance of object recognition, which best represents the VQA task, under various typical typographic factors, as illustrated in Fig. 6 (a). More detailed results with more various factors of other sub-tasks are shown in Appendix Sec. 0.A. Regarding Font Size and Opacity, Fig. 6 (a) shows a conspicuous positive correlation between the visibility of typographic text and the efficacy of typographic attacks. Furthermore, less visible typographic text still results in a significant reduction in models’ accuracy. typographic text with a font size of 6 pixels and an opacity of 20% can still cause an 11% and 16.6% drop in accuracy, respectively. Concerning Spatial Positioning and Font Color, Fig. 6 (a) shows that typographic text, regardless of its color or position, can significantly impact the accuracy of models. The phenomenon of typographic attacks to remain effective despite variations in color and location may be utilized to craft images laden with typographic text that, while seemingly ordinary, contain inconspicuous errors that can mislead LVLMs. This can be achieved by selecting a font color that merges with the background, making it hard to spot, and positioning it in an area that tends not to draw the eye.
According to the exploration results of typographic factors in Fig. 6 and results in Appendix Sec. 0.A, we fix the final typographic factors as {15px, 100%, R2C2 and White}. Using this setting, the sub-datasets, namely TypoD-B and TypoD-L, in the Factor Fixing stage of our TypoD are introduced. Next, we validate the performance degradation of LVLMs on Obj, Vis, Enu, and Rea in Typo-D and Typo-L. In Typo-B, compared to the performance of LLaVA-v1.5 and InstructBLIP in WTypo-B, there would occur and performances drop. In Typo-L, the performances drop compared to WTypo-L are and respectively. We observe that typographic attacks indeed significantly degrade the performance of LVLMs. However, there are differences in the robustness of different LVLMs against this threat, with InstructBLIP outperforming LLaVA-v1.5. For a more detailed exploration, we utilized LLaVA-v1.5 in our exploratory phase of LVLMs.
@@
Exploring Discoveries: In order to further validate the discoveries in Fig. 4, we initially examine more image instances in Appendix Sec. 0.B to observe the similarity score in vision-language modal comparison and verify the distract effects of visual attention in Grad-CAM images. In addition, to assess the impact of changes in input information of the text encoder of CLIP on the severity of typographic threats in the zero-shot classification task, we randomly select 500 images from ImageNet and create their typographic versions by printing a random class as a typographic error in images. Hence we get a set of normal images and a set of typographic images. Then we craft image-matching texts for these images using Set 1 and Set 2 templates. Set 1: [an image of {label}, an image of {typo}]. Set 2: [an image of {label} with a word {typo} written on top of it, an image of {typo} with a word {typo} written on top of it, an image of {label}, an image of {typo}]. Through separate testing on CLIP-Base (CLIP-B) and CLIP-Large (CLIP-L), according to the Fig. 6, adopting the more informative Set 2 as templates would result in the enhancements of and compared to the less informative Set 1 in the zero-shot classification task.
For validating the discoveries in Fig. 5, we need to augment prompts with additional semantics related to "describing image content" to obtain new responses in text language modality. The text response generated by this new prompt, according to Discovery 2, on the one hand, could mitigate the visual attention already attracted by the typographic text due to the increased information in the text modality. Also according to Discovery 3, on the other hand, the additional information in the text modality can also match attention with the prompt, further directing overall attention favorably towards the correct response. we specifically implement the semantics of the above prompts by 1. Directly adding requirements for describing images to the prompt; 2. Using multi-step prompts to generate more text language responses. Exemplary instructed prompts are illustrated following.
@@
Compared to the GAP of for LLaVA-v1.5 in TypoD-B as shown in Table 2, the three prompts we test in Table 3 result in reductions in GAP to , , and , respectively. This indicates that adding semantics related to image description in prompts is indeed effective, but there may be a threshold, which may arise from the inherent gap in querying between multi-modal information. Additionally, by comparing the GAP of , , and for Prompt 2.1, 2.2, and 2.3, it further demonstrates that through the incremental increase of semantics describing image content, there is a slight stair-step suppression effect on typographic threats. More results, such as evaluation on various prompts and models, are presented in Appendix Sec. 0.B.
Ablation Study: Except for the above four multi-modal sub-tasks, in Appendix Sec. 0.C we also explore the impact of typography on images containing texts. Additionally, we also tested a more intuitive approach to prompt semantic modification for typographic attacks. As presented in Fig. 6 (c), a further GAP decrease could be attained by prompts instructed to ignore typography modified from the above informative prompts.
@@ 6 Conclusion
In this paper, we conduct verification of typographic attacks on leading LVLMs and confirm their prevalent existence. We presented an extensive Typographic Dataset to delve deeper into this susceptibility, evaluating typographic risks across different multi-modal tasks and typographic factors.
Appendix 0.A Evaluation on TypoD
0.A.1 Details of Evaluated LVLMs
@@
0.A.2 More Quantitative Results
@@
Appendix 0.B Exploring Discoveries
0.B.1 CLIP with Informative Text Encoder Inputs
CLIP could almost avoid typographic attacks in zero-shot classification if more informative prompts are provided, as demonstrated in Fig. 4. In this subsection, we provide more examples to illustrate above discovery as presented in Fig. 7. By providing additional information to the prompt text, such as including more options like "an image of a {stingray, tape player} with the word ’tape player’ written on it, CLIP shifts its attention area and correctly distinguishes between the visual contents and the added typographic texts under a typographic attack. This demonstrates that CLIP itself is not flawed and its vision encoder can accurately extract complete information. The crucial issue lies in the varying levels of information contained in the text options provided for the text encoder.
0.B.2 LLaVA with Informative Prompts
0.B.3 Performance of Different Informative Prompts in LLaVA-v1.6 and InstructBLIP
For further validating the discoveries in Fig. 5, we evaluate informative prompts in Table 3 on LLaVA-v1.6 and InstructBLIP, as shown in Table 6 and Table 7. Compared to the GAP of for LLaVA-v1.6 in TypoD-B, the five prompts we test result in narrower GAPs of , , , and , respectively. This indicates that adding semantics related to image description in prompts is indeed effective.
@@ 0.B.2 LLaVA with Informative Prompts
Obj | 98.4 | 49.6 | 48.8 | 97.6 | 77.3 | 20.2 | 98.1 | 78.9 | 19.2 | 97.6 | 81.5 | 16.0 | 97.0 | 84.3 | 12.6 |
Vis | 92.3 | 82.0 | 10.2 | 97.9 | 94.8 | 3.1 | 97.9 | 95.8 | 2.0 | 96.9 | 95.3 | 1.5 | 97.4 | 95.3 | 2.0 |
Enu | 76.0 | 50.7 | 25.2 | 91.8 | 70.2 | 21.5 | 91.8 | 71.3 | 20.5 | 91.5 | 76.5 | 15.0 | 92.8 | 77.6 | 15.2 |
Rea | 90.3 | 44.8 | 45.4 | 89.5 | 60.0 | 29.4 | 86.2 | 63.6 | 22.5 | 86.1 | 67.6 | 18.4 | 83.2 | 67.6 | 15.5 |
@@
Appendix 0.C Ablation Study
0.C.1 Prompts Instructed to Ignore Typographic Texts
Additionally, we test a more intuitive approach to prompt semantic modification for typographic attacks on LLaVA-v1.5. We add the additional words "unaffected by typographic texts." to each Prompts mentioned in Table 3 respectively. Then we get new Prompt 1-1, Prompt 2-1 and Prompt 3-1 which are instructed to ignore typographic texts. As shown in Table 8, compared to GAPs of unmodified prompts in Table 3 which are , and respectively, their instructed versions of ignoring typographic texts result in narrower GAPs of , , . This suggests that incorporating instructions to overlook typographic texts within prompts can genuinely enhance performance.
0.C.2 Detailed Results of Typogrpahic Factors in differernt sub-tasks
In order to further evaluate how typography affect the performance of LVLMs. We evaluate LLaVA-v1.5 and InstructBLIP in TypoD-B with more various typographic factors in four sub-tasks. The results of object recognition, visual attribute detection, enumeration, and commonsense reasoning are presented in Table 9, Table 10, Table 11 and Table 12.
Additionally, we also craft a typographic arithmetic dataset. In the arithmetic computation, LVLMs need to comprehend arithmetic expressions embedded in images. Here, we focus on relatively easy arithmetic problems like addition, subtraction, multiplication, and division. We randomly selected two numbers and chose an arithmetic operation to compose an arithmetic problem. The correct answer is assigned as the ground truth, and a random number is selected as a typo. Then both the arithmetic problem and the typographic text are printed into the image. The distractibility of LVLMs with different typographic factors in the arithmetic computation task is shown in Table 13.